OpenStudy (anonymous):

How does the chain rule apply to \[ \frac{d}{dx} ln(y*x) \] with y a constant

OpenStudy (ksaimouli):



OpenStudy (ksaimouli):

y some number so derivative x is 1

OpenStudy (anonymous):

wut

OpenStudy (anonymous):

I'm not asking the answer I'm asking how the chain rule applies to this problem

OpenStudy (ksaimouli):

what is derivative of lnx

OpenStudy (anonymous):

1/x

OpenStudy (ksaimouli):

so what about lnIxyI

OpenStudy (anonymous):

1/xy?

OpenStudy (ksaimouli):

yes but u did not take derivative of xy because u took derivative of ln

OpenStudy (ksaimouli):

next step is to take derivative of xy but y is constant

OpenStudy (anonymous):

Why do we split this up? How can I use chain rule here? Because I don't see any function coposition going on...

OpenStudy (anonymous):

function composition*

OpenStudy (anonymous):

I don't need to know the next step I need to know WHY

OpenStudy (abb0t):

The derivative of ln and then the derivative of (xy), using product rule.

OpenStudy (abb0t):

Chain rule states: \[\frac{ df }{ dx } = \frac{ df }{ dg }(\frac{ dg }{ dx })\]